If bob has $1,147.80 and it is in a savings account for 42 months with an interest rat of 4.4% whats his new total?
Answer by stanbon(75887)

If bob has $1,147.80 and it is in a savings account for 42 months with an interest rat of 4.4% whats his new total?
------
A(t) = P(1 + rt)
-----
A(42/12) = 1147.80(1+0.044*(42/12))
----
A(42/12) = $1324.56
========================
Cheers,
Stan H.